write the slope intercept equation for (-4,-1) and (1,9)
Answer by rfer(16322)

slope
m=(9--1)/(1--4)
m=10/5
m=2
---------------
y=mx+b
9=2*1+b
9=2+b
b=4.5
----------------
y=2x+4.5